# Impact of Forecast Stability on Navigational Contrail Avoidance

### v1

@@ Abstract
Mitigating contrail-induced warming by re-routing flights around contrail-forming regions requires accurate and stable forecasts of the state of the upper troposphere and lower stratosphere. Forecast stability (i.e., consistency between forecast cycles with different lead times) is particularly important for “pre-tactical” contrail avoidance strategies that adjust routes based on forecasts with lead times as long as 24-48 hours. However, no study to date has systematically quantified the degree to which forecast stability limits the effectiveness of pre-tactical avoidance. This study addresses this gap by comparing contrail forecasts generated using ECMWF HRES weather forecasts with lead times up to 48 hours to contrail hindcasts generated based on ECMWF ERA5 reanalysis. An analysis of forecast errors shows low pointwise consistency between persistent-contrail-forming regions in forecasts and reanalysis, with pointwise error rates similar to those found in previous comparisons of contrail-forming regions in reanalysis and reality. However, we also show that spatial errors in the locations of contrail-forming regions are relatively small, both when forecasts are compared to reanalysis and when reanalysis is compared to in-situ measurements. Finally, we show that designing a trajectory optimizer to take advantage of relatively small spatial errors allows flight trajectory optimizations based on contrail forecasts to reduce contrail climate forcing evaluated based on reanalysis by 80-90% at the 8-24 hour lead times most relevant to flight planning, with fuel penalties under 0.4%. Our results show that forecasts with lead times relevant to flight planning are stable enough to be used for pre-tactical contrail avoidance.
@@ 1 Introduction
Persistent contrails and contrail-induced cirrus clouds are responsible for about half of the current anthropogenic climate forcing from aviation [1]. Contrails are ice clouds that form when jet engine exhaust, which consists of combustion byproducts including particulate matter and water vapor, satisfies the Schmidt-Appleman criterion [2]. If the Schmidt-Appleman criterion (SAC) is satisfied, the exhaust plume becomes supersaturated with respect to water during mixing with cold ambient air, allowing water vapor to condense onto the surface of particles found in the exhaust plume and subsequently freeze to form ice crystals [3, 4]. If the ambient atmosphere is supersaturated with respect to ice (i.e., relative humidity with respect to ice, RHi, is above 100%), then the ice particles will persist and the contrail will spread, over time becoming indistinguishable from natural cirrus clouds [5, 6, 7]. Estimates of current effective radiative forcing from persistent contrails and contrail cirrus are comparable to, and in some cases higher than, estimates of radiative forcing from historical aviation CO2 emissions [8, 9, 10, 1].
[…]
@@
A third approach to mitigating the climate impact of contrails is through navigational avoidance of regions where warming contrails can form [22, 23, 24, 16, 25]. The simplest form of navigational avoidance—avoidance of ice super-saturated regions (ISSRs) that support persistent contrails—may be feasible, but could require avoiding a prohibitively large volume of airspace. However, avoiding all ISSRs may be unnecessary, as some persistent contrails are relatively short-lived and some have a net cooling effect [26]. Accordingly, several studies have proposed using forecasts of contrail lifetime and radiative forcing (generated e.g. based on the contrail cirrus prediction model, or CoCiP) to avoid only the subset of ISSRs where particularly long-lived or strongly-warming contrails form [27, 26, 28].
@@
This paper focuses specifically on “pre-tactical” avoidance, where contrail-aware flight plans are generated at the time of flight planning. This approach differs from “strategic” avoidance, which relies on climatological properties of contrails (e.g., by reducing flight distance at night when contrails are almost always net warming), and “tactical” avoidance, which relies on requesting mid-flight trajectory changes to avoid contrail-forming regions. For commercial airlines with regularly-scheduled flights, optimized flight trajectories are typically generated by flight planning software within 24 hours of the flight’s departure and then manually adjusted by flight crews to account for en-route hazards and other operational issues such as congestion. Depending on the demands of the operator, flight trajectories may be re-optimized within a few hours departure using the most up-to-date weather forecasts. Flight planning at lead times exceeding 24 hours may occur in special situations, such as planning for cargo capacity on long-haul routes, or due to delays in forecast delivery.
@@
Because forecasts are free-running and include errors that grow with increasing lead time, whereas reanalyses are continuously constrained by assimilated observations, optimizing trajectories using forecasts and evaluating reductions in contrail warming using reanalysis provides a way to isolate the impact of forecast error growth. We emphasize, however, that this approach does not provide direct insight into the impact of forecast accuracy (i.e., differences between forecasts and reality) on the effectiveness of navigational avoidance because reanalysis is not a perfect source of meteorological truth.
@@ 2.1 Meteorology data
We used the European Center for Medium-Range Weather Forecasting (ECMWF) IFS HRES as our source of weather forecasts and the ECMWF ERA5 HRES reanalysis [34] as our source of reanalysis data [^1] Note that this manuscript uses “forecast” or “HRES” to refer to ECMWF IFS HRES forecasts and “reanalysis“ or “ERA5” to refer to ECMWF ERA5 HRES reanalysis.. We retrieved forecast and reanalysis fields on model levels at 0.25 degree horizontal resolution. Unless otherwise noted, forecast fields were retrieved from the ECMWF Operational Archive and reanalysis fields from the Copernicus Climate Data Store. In all cases, we preprocessed forecast and reanalysis data before further use by interpolating fields onto 30 pressure levels roughly equally spaced between FL200 and FL500 and by applying the quantile-matching humidity scaling described by [35] to forecast and reanalysis specific humidity.
@@ 2.2 Contrail forecasts
We generated forecasts of contrail warming using the CoCiP grid model [28] with HRES forecasts as input. The CoCiP grid model is an approximation of the original “trajectory” version of CoCiP [27], a Lagrangian Gaussian plume model that estimates contrail properties for complete flight trajectories. Rather than estimating properties for full flight trajectories, CoCiP grid simulates the evolution of infinitesimal contrail segments initialized on a 4D spatiotemporal grid to produce a gridded estimate of contrail energy forcing (EF) produced by flights through different volumes of airspace. This requires making simplifying assumptions about aircraft performance and contrail orientation (see [28] for details), but allows CoCiP grid to efficiently compute global forecasts of contrail EF in a format that can be easily integrated into commercial flight planning software.
@@ 2.3 Forecast error analysis
We quantified the frequency of pointwise errors in contrail forecasts using the equitable threat score (ETS) metric [29, 31]. For a set of predictions of a binary variable with $a$ true positives, $b$ false positives, $c$ false negatives, and $d$ true negatives, the ETS is
@@
where $r=(a+b)(a+c)/(a+b+c+d)$ is the expected number of true positives produced by a random model. The ETS for a random model is 0 and the ETS for a perfect model is 1. With a low base rate (far fewer positives than negatives), the ETS for a model with precision and recall of 0.5 is $1/3$. To ensure that ETS scores reflected the frequency of errors at locations relevant to flight planning, we computed ETS scores from predictions interpolated to flight waypoints.
Pointwise measures of contrail forecast accuracy are vulnerable to the so-called “double-penalty” issue, widely discussed in the literature on evaluating high-resolution forecasts (e.g., [36]). Because contrail forecasts predict localized features, relatively small displacement errors in the predicted locations of contrail-forming regions can produce very frequent pointwise errors. To address this limitation, we also quantified the proximity of forecasted contrail-forming regions to contrail-forming regions found in reanalysis based on a “proximity distribution”. To compute proximity distributions, we took sets of flight trajectories, calculated the flight time from each flight waypoint to the nearest forecasted contrail-forming region along the same flight trajectory (assigning a flight time of infinity for waypoints on flight trajectories for which no contrails were forecasted), and computed probability distributions from that set of flight times. The proximity distribution for flight waypoints in true contrail-forming regions provides a measure of their proximity to contrail-forming regions in forecasts.
@@ 2.4 Statistical Methods
Unless otherwise stated, statistics presented as a function of forecast lead time were computed by binning waypoints based on a single nominal lead time assigned to all waypoints in a flight. Nominal lead times were computed as the difference between the forecast initialization time and the time of the flight’s first waypoint. Error bars for binned statistics correspond to 95% confidence intervals obtained by applying a bias-corrected accelerated (BCa) bootstrap analysis [37]. For each statistic, the bootstrap samples were obtained by resampling (with replacement and before binning) from the set of all flights considered in this study (i.e., they are resampled per flight and not per day or per waypoint). Each statistic was resampled a total of 10,000 times to obtain a bootstrap distribution. The confidence intervals were chosen from the bootstrap distribution according to the BCa method originally described by [37], which accounts for both bias and skewness in the bootstrap distribution.
@@ 3 Optimizer design
The optimizer performs a breadth-first Dykstra-like search across the 2D grid to find the vertical profile that minimizes the objective function
@@
Fuel burn was estimated using the Poll-Schumann aircraft performance model [38, 39], as implemented in [40]. The Poll-Schumann model also provides estimates of thrust limitations and flight envelope restrictions (i.e., minimum and maximum permitted Mach numbers), and we ensured that all optimized trajectories satisfied these constraints. We estimated the initial mass of each flight based on an assumed load factor and reserve fuel requirements following [6], and assumed a default engine type based on the mapping from aircraft type provided by [6]. We set the cost index to a fixed value of 70 kg/min, effectively assuming that operating an aircraft for one minute costs the equivalent of burning 70 kg of fuel. In practice, this value varies highly between operators and operational conditions. The contrail cost index was set to 0 for cost-optimal trajectories. For contrail-optimal trajectories, the default contrail cost index was set to $6.7\times 10^{-11}\,\mathrm{kg/J}$, which converts contrail EF to a CO2-equivalent fuel burn based on an absolute global warming potential over 100 years (AGWP100) [41] [^2] Note that there is not a straightforward relationship between the contrail cost index and the actual fuel penalty paid to fly contrail-optimal rather than cost-optimal trajectories. Using a cost index based on AGWP100 does not imply that the fuel penalty will be equal and opposite to the reduction in contrail warming on AGWP100 terms. The cost index can be tuned to produce different tradeoffs between contrail impacts and costs; a further exploration of this is presented in A.. Contrail EF at each waypoint was computed based on a 4D interpolation of gridded CoCiP forecasts for the appropriate aircraft-engine class. Other costs, such as overflight charges or engine cycle costs, were not considered. Example cost- and contrail-optimal trajectories are shown in Figure 1.
Figure 1: Two example optimized flight trajectories, showing for each a cost-optimal trajectory (dashed line) and a contrail-optimal trajectory (solid line). The trajectory on the left-hand side reduces contrail EF by $1.2\times 10^{15}\,\mathrm{J}$ ($340,000\,\mathrm{kg}$ CO2eAGWP100), with an additional fuel burn of $2,233\,\mathrm{kg}$ ($+4.7\%$). The trajectory on the right-hand side saves $6.3\times 10^{13}\,\mathrm{J}$ ($18,000\,\mathrm{kg}$ CO2eAGWP100) of contrail EF, with a fuel penalty of $200\,\mathrm{kg}$ ($+2.4\%$).

### v3

@@ Abstract
Mitigating contrail-induced warming by re-routing flights around contrail-forming regions requires accurate and stable forecasts of the state of the upper troposphere and lower stratosphere. Forecast stability (i.e., consistency between forecast cycles with different lead times) is particularly important for “pre-tactical” contrail avoidance strategies that adjust routes based on forecasts with lead times as long as 24-48 hours. However, no study to date has systematically quantified the degree to which forecast stability limits the effectiveness of pre-tactical avoidance. This study addresses this gap by comparing contrail forecasts generated using ECMWF HRES weather forecasts with lead times up to 48 hours to contrail hindcasts generated based on ECMWF ERA5 reanalysis. An analysis of forecast errors show low pointwise consistency between persistent-contrail-forming regions in forecasts and reanalysis, with pointwise error rates similar to those found in previous comparisons of contrail-forming regions in reanalysis and reality. However, we also show that spatial errors in the locations of contrail-forming regions are relatively small, both when forecasts are compared to reanalysis and when reanalysis is compared to in-situ measurements. Finally, we show that designing a trajectory optimizer to take advantage of relatively small spatial errors allows flight trajectory optimizations based on contrail forecasts to reduce contrail climate forcing evaluated based on reanalysis by 80-90% at the 8-24 hour lead times most relevant to flight planning, with fuel penalties under 0.4%. Our results show that forecasts with lead times relevant to flight planning are stable enough to be used for pre-tactical contrail avoidance.
@@ 1 Introduction
Persistent contrails and contrail-induced cirrus clouds are responsible for about half of the current anthropogenic climate forcing from aviation [1]. Contrails are ice clouds that can form when the jet engine exhaust plume, which consists of combustion byproducts including particulate matter and water vapor, mixes with cold ambient air and becomes supersaturated with respect to water. This occurs when the Schmidt-Appleman criterion (SAC) is satisfied (i.e., when the ambient temperature falls below the SAC threshold temperature), causing water vapor to condense onto the surface of particles found in the exhaust plume and subsequently freeze to form ice crystals [2, 3, 4]. If the ambient atmosphere is supersaturated with respect to ice (i.e., relative humidity with respect to ice, RHi, is above 100%), then the ice particles will persist and the contrail will spread, over time becoming indistinguishable from natural cirrus clouds [5, 6, 7]. Estimates of current effective radiative forcing from persistent contrails and contrail cirrus are comparable to, and in some cases higher than, estimates of radiative forcing from historical aviation CO2 emissions [1, 8, 9, 10].
There are several proposed solutions to mitigate the climate impacts of contrails. Some studies have suggested that contrails formed by flights using sustainable aviation fuel (SAF) may be optically thinner, shorter-lived, and less strongly warming than contrails formed by flights using conventional jet fuels [11, 12, 13]. However, recent studies have identified several challenges in using SAF to reduce contrail climate forcing. The costs of SAF-based contrail abatement ($\mathrm{EUR}\,14$-61/tCO2e) are likely significantly higher than the costs of navigational avoidance ($\mathrm{EUR}\,1-$3/tCO2e) after accounting SAF supply chain costs [14]. Additionally, SAF availability remains low ($\sim 0.2\%$ of the global annual aviation fuel consumption as of 2023) with slow growth rates [15, 16]. Thus, while SAF may play an important role in decarbonizing aviation, its potential for near-term contrail mitigation is limited.
Another proposed mitigation strategy is the use of lean-burn combustors as an engine option [17]. Lean-burn combustors emit up to five orders of magnitude less non-volatile particulate matter (nvPM) than conventional engines [18] and could, in principle, produce optically-thinner and shorter-lived contrails if water droplets nucleated exclusively on nvPM. However, recent in-flight measurements reported initial apparent contrail ice emissions index on the order of $10^{14}$ to $10^{15}$ kg-1, comparable to those from conventional engines, despite the nvPM number emissions index being only around $10^{11}$ kg-1. This discrepancy may be explained by the presence of volatile particulate matter (vPM, i.e., a mixture of engine lubrication oil droplets, sulfuric acid, and organic compounds), which likely serve as condensation nuclei when nvPM is scarce [4, 19, 20, 21].
@@
A third approach to mitigating the climate impact of contrails is through navigational avoidance of regions where warming contrails can form [16, 22, 23, 24, 25]. The simplest form of navigational avoidance—avoidance of ice super-saturated regions (ISSRs) that support persistent contrails—may be feasible, but could require avoiding a prohibitively large volume of airspace, as quantified in [26]. However, significant contrail mitigation is still possible without avoiding all ISSRs, as around 80% of contrails do not persist beyond 5 minutes and around 30% of persistent contrails have a net cooling effect [6]. Accordingly, several studies have proposed using forecasts of contrail lifetime and radiative forcing (generated e.g. based on the contrail cirrus prediction model, or CoCiP) to avoid only the subset of ISSRs where particularly long-lived or strongly-warming contrails form [27, 28, 26].
Navigational avoidance strategies will likely cause some amount of increased fuel consumption and a proportional increase CO2 emissions [16]. Due to the differing nature of the lifetime and uncertainties associated with each of these forcing agents, quantifying overall climate costs and benefits associated with a contrail mitigation scheme is a complex topic and outside of the scope of this work. However, we note that recent studies have suggested that navigational avoidance of contrails is highly likely to have a positive climate outcome despite uncertainties about the magnitude of contrail climate impacts and the most appropriate choice of impact metric [29, 30, 31].
@@
This paper focuses specifically on “pre-tactical” avoidance, where contrail-aware flight plans are generated at the time of flight planning. This approach differs from “strategic” avoidance, which relies on climatological properties of contrails (e.g., by reducing flight distance at night when contrails are almost always net warming), and “tactical” avoidance, which relies on requesting mid-flight trajectory changes to avoid contrail-forming regions. For commercial airlines with regularly-scheduled flights, optimized flight trajectories are typically generated by flight planning software within 24 hours of the flight’s departure. Routes are optimized according to a set of criteria that will vary by operator, but will include some combination of minimizing fuel, time, and other operational costs. Optimized trajectories are manually adjusted by flight crews and air traffic controllers to account for en-route hazards and other operational issues such as airspace congestion. Depending on the demands of the operator, flight trajectories may be re-optimized within a few hours departure using the most up-to-date weather forecasts. Flight planning at lead times exceeding 24 hours may occur in special situations, such as planning for cargo capacity on long-haul routes, or due to delays in forecast delivery.
@@
Weather forecasts are initialized from meteorological state estimates (“analyses”) produced by data assimilation systems that combine model outputs with observations. Because forecasts are free-running, they include errors that grow with increasing lead time. In contrast, reanalyses are continuously constrained by assimilated observations, and optimizing trajectories using forecasts and evaluating reductions in contrail warming using reanalysis therefore provides a way to examine the impact of forecast error growth. (It also accounts for the impact of errors in analyses, which are required to be available in near-real-time, relative to reanalysis, which is produced retrospectively and can assimilate a larger set of observations.) We emphasize, however, that this approach does not provide direct insight into the impact of forecast accuracy (i.e., differences between forecasts and reality) on the effectiveness of navigational avoidance because reanalysis is not a perfect source of meteorological truth.
@@ 2.1 Meteorology data
We used the European Center for Medium-Range Weather Forecasts (ECMWF) IFS HRES as our source of weather forecasts and the ECMWF ERA5 HRES reanalysis [37] as our source of reanalysis data [^1] Note that this manuscript uses “forecast” or “HRES” to refer to ECMWF IFS HRES forecasts and “reanalysis“ or “ERA5” to refer to ECMWF ERA5 HRES reanalysis.. We retrieved forecast and reanalysis fields on model levels at 0.25 degree horizontal resolution. Unless otherwise noted, forecast fields were retrieved from the ECMWF Operational Archive and reanalysis fields from the Copernicus Climate Data Store. In all cases, we preprocessed forecast and reanalysis data before further use by interpolating fields onto 30 pressure levels roughly equally spaced between FL200 and FL500 and by applying the quantile-matching humidity scaling described by [38] to forecast and reanalysis specific humidity.
@@ 2.2 Contrail forecasts
We generated forecasts of contrail warming using the CoCiP grid model [26] with HRES forecasts as input. The CoCiP grid model is an approximation of the original “trajectory” version of CoCiP [28], a Lagrangian Gaussian plume model that estimates contrail properties for complete flight trajectories. Rather than estimating properties for full flight trajectories, CoCiP grid simulates the evolution of infinitesimal contrail segments initialized on a 4D spatiotemporal grid to produce a gridded estimate of contrail energy forcing (EF) produced by flights through different volumes of airspace. This requires making simplifying assumptions about aircraft performance and contrail orientation (see [26] for details), but allows CoCiP grid to efficiently compute global forecasts of contrail EF in a format that can be easily integrated into commercial flight planning software. The metric of per-waypoint EF is defined using the same conventions as the original formulation of CoCiP [28], which is obtained by integrated the radiative flux across the area and lifetime of contrail segment formed between a flight waypoint and its subsequent waypoint.
@@
Throughout this work, we use CoCiP grid to estimate contrail impacts when performing trajectory optimization. When possible, contrail impacts of an individual flight trajectory are computed using the higher-fidelity trajectory version of CoCiP, accounting for the correct aircraft type, and engine type. The trajectory version of CoCiP estimates fuel flow rates across the aircraft trajectory using an assumed load fact, providing more accurate estimates of nvPM and initial ice concentration numbers than the grid variant of CoCiP which relies on nominal engine efficiency values to estimate fuel consumption.
@@ 2.3 Forecast error analysis
We quantified the frequency of pointwise errors in contrail forecasts using the equitable threat score (ETS) metric [32, 34]. Unless otherwise noted, we use “pointwise” to denote metrics computed per flight waypoint. For a set of predictions of a binary variable with $a$ true positives, $b$ false positives, $c$ false negatives, and $d$ true negatives, the ETS is
@@
where $r=(a+b)(a+c)/(a+b+c+d)$ is the expected number of true positives produced by random matches. The ETS for a random model is 0 and the ETS for a perfect model is 1. With a low base rate (far fewer positives than negatives), the ETS for a model with precision and recall of 0.5 is $1/3$. To ensure that ETS scores reflected the frequency of errors at locations relevant to flight planning, we computed ETS scores from predictions interpolated to flight waypoints.
Pointwise measures of contrail forecast accuracy are vulnerable to the so-called “double-penalty” issue, widely discussed in the literature on evaluating high-resolution forecasts (e.g., [39]). Because contrail forecasts predict localized features, relatively small displacement errors in the predicted locations of contrail-forming regions can produce very frequent pointwise errors. To address this limitation, we also quantified the proximity of forecasted contrail-forming regions to contrail-forming regions found in reanalysis based on a “proximity distribution”. To compute proximity distributions, we took sets of flight trajectories, calculated the flight time from each flight waypoint to the nearest forecasted contrail-forming region along the same flight trajectory (assigning a flight time of infinity for waypoints on flight trajectories for which no contrails were forecasted), and computed probability distributions from that set of flight times. The proximity distribution for flight waypoints in true contrail-forming regions provides a measure of their proximity to contrail-forming regions in forecasts. We note that because flights follow approximately horizontal trajectories, this metric primarily captures horizontal rather than vertical proximity.
@@ 2.4 Statistical Methods
Unless otherwise stated, statistics presented as a function of forecast lead time were computed by binning waypoints based on a single nominal lead time assigned to all waypoints in a flight. Nominal lead times were computed as the difference between the forecast initialization time and the time of the flight’s first waypoint. Error bars for binned statistics correspond to 95% confidence intervals obtained by applying a bias-corrected accelerated (BCa) bootstrap analysis [40], which we now briefly describe. For each statistic, bootstrap samples are obtained by resampling (with replacement and before binning) from the set of all flights considered in this study (i.e., they are resampled per flight and not per day or per waypoint). Each statistic is resampled a total of 10,000 times. The empirical collection of resampled statistics is known as a bootstrap distribution. The confidence intervals are chosen from the tails of this bootstrap distribution. In the most basic form of bootstrap, the confidence interval based on quantiles of the bootstrap distribution. However, if the bootstrap distribution is biased or skewed, then this method will not produce an accurate estimate of the confidence intervals. In this case, they may be chosen according the BCa method originally described by [40], which accounts for both bias and skewness in the bootstrap distribution.
@@ 3 Optimizer design
The optimizer finds a vertical profile that minimizes the objective function
@@
This is accomplished through a breadth-first Dykstra-like search across the 2D grid described above. Precisely, the optimizer explores all feasible vertical trajectories (search paths) across the 2D grid in order of increasing total cost, and returns the lowest-cost trajectory. A more complete description of the optimization procedure is described in Appendix A6 of [26]. Fuel burn was estimated using the Poll-Schumann aircraft performance model [41, 42], as implemented in [43]. The Poll-Schumann model also provides estimates of thrust limitations and flight envelope restrictions (i.e., minimum and maximum permitted Mach numbers), and we ensured that all optimized trajectories satisfied these constraints. We estimated the initial mass of each flight based on an assumed load factor and reserve fuel requirements following [6], and assumed a default engine type based on the mapping from aircraft type provided by [6]. We set the cost index to a fixed value of 70 kg/min, effectively assuming that operating an aircraft for one minute costs the equivalent of burning 70 kg of fuel. In practice, this value varies highly between operators and operational conditions. The contrail cost index was set to 0 for cost-optimal trajectories. For contrail-optimal trajectories, the default contrail cost index was set to $6.7\times 10^{-11}\,\mathrm{kg/J}$, which converts contrail EF to a CO2-equivalent fuel burn based on an absolute global warming potential over 100 years (AGWP100) [44] [^2] Note that there is not a straightforward relationship between the contrail cost index and the actual fuel penalty paid to fly contrail-optimal rather than cost-optimal trajectories. Using a cost index based on AGWP100 does not imply that the fuel penalty will be equal and opposite to the reduction in contrail warming on AGWP100 terms. The cost index can be tuned to produce different tradeoffs between contrail impacts and costs; a further exploration of this is presented in A.. Contrail EF at each waypoint was computed based on a 4D interpolation of gridded CoCiP forecasts for the appropriate aircraft-engine class. Other costs, such as overflight charges or engine cycle costs, were not considered. Example cost- and contrail-optimal trajectories are shown in Figure 1. It is worth noting that the contrail-optimal trajectories depicted in Figure 1 do not fully avoid contrail-forming regions, but instead find trajectories that balance contrail formation with excess fuel and time costs.
Figure 1: Two example optimized flight trajectories, showing for each a cost-optimal trajectory (dashed line) and a contrail-optimal trajectory (solid line). The trajectory on the left-hand side reduces contrail EF by $1.2\times 10^{15}\,\mathrm{J}$ ($340,000\,\mathrm{kg}$ CO2eAGWP100), with an additional fuel burn of $2,233\,\mathrm{kg}$ ($4.7\%$). The trajectory on the right-hand side saves $6.3\times 10^{13}\,\mathrm{J}$ ($18,000\,\mathrm{kg}$ CO2eAGWP100) of contrail EF, with a fuel penalty of $200\,\mathrm{kg}$ ($2.4\%$).